OpenStudy (anonymous):

solve the system of equations. I genuinely have no clue how to do this so if someone could help explain it, that would be fantastic! 1) x+y-z=-1 x+y+z=3 3x-2y-z=4 2)x+y=5 3x+z=2 4y-z=8 3) 2x-y+z=-1 x-y+z=1 x-2y+z=2 4) x-z=5 y+3z=12 2x+y=7

OpenStudy (anonymous):

Well these are actually really fun. Let's look at number 1. What equation there seems like the simplest to solve for a variable? (Which one will require the least amount of step/or is easiest for you) to solve for one of the variables in it.

OpenStudy (anonymous):

Ill help with number 2 since @Cosmichaotic will probably help you with number 1.

OpenStudy (anonymous):

Haha, nice.

OpenStudy (anonymous):

we can both help :D

OpenStudy (anonymous):

Of course, I prefer that. I make a lot of mistakes and need supervision.

OpenStudy (anonymous):

ok, for number 2. you want to end up finding what values for x, y, and z make each equation true. so what you probably want to do is focus first on solving for z in the second equation

OpenStudy (anonymous):

@cosmichaotic Well the last one has more numbers in it so it seems fairly straightforward I guess

OpenStudy (anonymous):

to do that, subtract your 3x from both sides so you have z = -3x +2

OpenStudy (anonymous):

next, look at your equation 3 for number 2 and plug in (-3x+2) for z

OpenStudy (anonymous):

Alright so it's your choice, do you want to solve for x, y or z?

OpenStudy (anonymous):

More numbers means more operations, let's choose the first equation since it seems simple enough to isolate one of the variables.

OpenStudy (anonymous):

what you should be looking at now is something such as this \[4y - (-3x + 2) = 8\]

OpenStudy (anonymous):

with all your negatives multiplied out you should end up with this: \[4y + 3x - 2 = 8\]

OpenStudy (anonymous):

@cosmichaotic okay, that makes sense... So we just start to solve for x?

OpenStudy (anonymous):

@technosoul And then I just solve for the next variable and plug it in to the next equation?

OpenStudy (anonymous):

pretty much, yes

OpenStudy (anonymous):

@technosoul Thank you so much! You have no idea how much it helped

OpenStudy (anonymous):

sure thing. just so you know though, number two is going to have a slightly confusing answer and i can explain that after you find your answer

OpenStudy (anonymous):

ill start on number 3 though, for now.

OpenStudy (anonymous):

Absolutely, sure, solve for x. We will get: x+y-z=-1 -> x = -1 -y + z So now that we have x, (x = -1 -y + z), we can plug that into the x in the next equation. Plugging in what we got for x: (x)+y+z=3 -> -1 -y + z + y + z = 3

OpenStudy (anonymous):

Now we just have two variables to worry about. First, let's simplify it by combining like terms. -1 -y + z + y + z = 3 Adding the y's and z's, where -y + y = 0 (We won't have to worry about y's) and z + z = 2z We have, -1 +0y + 2z = 3 or -1 + 2z = 3 Adding 1 to both sides to combine the constants, -1 + 2z = 3 -> 2z = 4 And we simplify this by dividing by 2 on both sides, 2z = 4 -> z = 2

OpenStudy (anonymous):

I know that's a long explanation, but read through it and let me know if you have any questions about how we came to: x = -1 -y + z z = 2 So far.

OpenStudy (anonymous):

alright, number 3. so start by solving for a variable (i hope its ok i chose to solve for z first). Your best option to solving for z is your first and your last equations and ill explain why. The first equation has a value of z that will cancel out the second equation's value of y. and therefore solving for your x. The last equation has a value of z that will cancel out the second problems value of x and therefore solving for your y.

OpenStudy (anonymous):

i promise it will make more sense when i write it out.

OpenStudy (anonymous):

so first solve for z for your first equation. \[2x - y +z=-1\] then you add y to both sides to get \[2x + z = y-1\] after that subtract the 2x from both sides and you'll get \[z = -2x +y-1\] make sense so far?

OpenStudy (anonymous):

I think that to solve it from there, your second equation is the easiest to use so plug in what you have for z in, in the second equation. You should get: \[x-y+(-2x+y-1)=1\]

OpenStudy (anonymous):

as i mentioned earlier, this will now solve for your x because your y's will now cancel out

OpenStudy (anonymous):

once you get that, tell me, i can check to make sure you got what i did